E304 Calculate E(0.87] where X is a Binomial random variable with the parameters (n=7, p=0.33). Answer: CHECK SE
Answer #1

E(0.87 . Y -> Binomial Rov %3A ।। - 20.33 P(X) = na no pa (1-2) 1.02,314,5,6,7 F(23) TOP(x=.) . M . ।। 2087) x=0 (0.33) 01-0.As (0.33 (0677 = 0.03689 RIXES) PGXD) ty (0.33) (6.67) P(x = 1) - 7 (933) (06230 0.00606 - Or66043 2 Х 3 P(X) 0.06061 0.20895